OpenStudy (anonymous):

In a circle, a 90 degree sector has area 16 pi ft^2. What is the radius of the circle?

OpenStudy (anonymous):

16 pi = pi r^2/4 find r

OpenStudy (anonymous):

90 deg is one fourth of full circle.. 16π = πr^2/4 4 * 16π ------- = r^2 π 64 = r^2 8 = r so radius = 8 ft